OpenStudy (angelwings996):

Please help!? Write an equation for a horizontal translation of y = 2/x. Then write an equation for a vertical translation of y = 2/x. Identify the horizontal and vertical asymptotes of the graph of each function.

OpenStudy (whpalmer4):

If you want to shift the equation right or left, you just add or subtract from the argument to the function. For example, a parabola \(y=x^2\) opens up, has its vertex at (0,0). If we change the argument of \(y = f(x) = x^2\) to \(x-2\), then it becomes \[y = f(x-2) = (x-2)^2\] and if you plot a few test points between x =-2 and x = 2 you'll see that we've just shifted the parabola 2 units along the x-axis. What originally happened at x =0 now happens at x = 2, and so on. For a vertical translation, that's just adding or subtracting from the result of the function. Going back to our parabola, if we instead wanted to shift it up 2 units, we would write\[y = f(x)+2 = x^2 +2\] and again plotting a few points will show that the parabola looks identical, except shifted upward 2 units. In particular, the vertex now comes at (0,2) because \[y = x^2+2 = 0^2 + 2\] So how would you translate your function in the two requested ways?